Compute Discounted Payback Period, Net Present Value and Profitability Index of a

project which costs Rs 50,000, cost of capital is 8% and is expected to generate

following cash inflows:Year Cash inflows(Rs)

1 12,000

2 15,000

3 25,000

4 28,000

5 20,000

6 22,000

On the basis of Net Present Value and Profitability Index, suggest if the project is to be

accepted or not.


1
Expert's answer
2021-04-06T05:44:05-0400

Discounted "CF = CF\/1.08^n."

Year CF Discounted CF

1 12,000 11,111.11

2 15,000 12,860.08

3 25,000 19,845.81

4 28,000 20,580.84

5 20,000 13,611.66

6 22,000 13,863.73

The Discounted Payback Period is approximately 3.3 years.

"NPV = -50,000 + 11,111.11 + 12,860.08 + 19,845.81 + 20,580.84 + 13,611.66 + 13,863.73 = 41,873.23."

Profitability Index = 41,873.23/50,000 = 0.837.

As NPV > 0 and PI is high enough, then the project is to be accepted.